OpenStudy (anonymous):

The equation for the circle below is x2 + y2 = 25. What is the length of the circle's radius?

myininaya (myininaya):

the radius is given you just have to know how to compare that equation to the standard form equation of a circle \[(x-h)^2+(y-k)^2=r^2\] we have here that h and k are both 0 so let's right this a little different since we know that \[x^2+y^2=r^2\] r represents radius

myininaya (myininaya):

comparing x^2+y^2=r^2 to x^2+y^2=25 what is 25?

OpenStudy (anonymous):

the radius ?

myininaya (myininaya):

r is the radius

myininaya (myininaya):

what is 25

OpenStudy (anonymous):

y=25

myininaya (myininaya):

look and tried to compare these two equations  it should be clear to you what 25 is 25 is in the place of?